OpenStudy (anonymous):

Someone helps me, please. (-8+4i)=(3+i)(-2+2i) It is true, but how to get the right hand side from the left hand side?

OpenStudy (tkhunny):

Essentially, you don't. It is kind of a factoring exercise, but really requires that you know the end from the beginning. Complex Numbers don't necessarily have unique factorizations. Look up "Gaussian Primes". You are asking, 12 = 3*4 -- How!!! Really, 12 = 2*6 is just as good. Or, 12 = 1*12. It's only unique if it's prime.

OpenStudy (mathmath333):

u can use "back solving".

OpenStudy (anonymous):

Don't get :) what is back solving?

OpenStudy (mathmath333):

u need to convert solution to the question.

OpenStudy (anonymous):

It's kind of magic, like \(-20-3\sqrt{13} =(1-\sqrt{13})(6+2\sqrt{13})\) There must exist some rule to get it.

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@iambatman @iGreen @IrishBoy123

OpenStudy (anonymous):

I don't solve it, just want to know how to go to the right hand side quickly like that. :)

OpenStudy (tkhunny):

"back solve" means to get the LHS from the RHS, or what I described as "know the end from the beginning." It's not necessarily unique. Stop trying to make it unique.